# NUBO: A Transparent Python Package for Bayesian Optimization

### v2

@@ 5 Conclusion
This article introduces NUBO, a Python package for Bayesian optimization to optimize expensive-to-evaluate black-box functions, for example, computer simulators and physical experiments. The main objective of NUBO is to make Bayesian optimization accessible to researchers from all disciplines by providing a transparent and user-friendly implementation.
NUBO includes five sub-modules that implement Gaussian processes, acquisition functions, optimizers, test functions, and utilities. These modules provide all necessities for sequential single-point, parallel multi-point, and asynchronous optimization of expensive-to-evaluate black-box functions for bounded, constrained, and mixed (discrete and continuous) input parameter spaces. We have introduced and explained each of these functionalities with individual code snippets and illustrated NUBO’s general workflow using a detailed case study that takes a hypothetical six-dimensional expensive-to-evaluate black-box function and approximates its global optimum with a parallel multi-point Bayesian optimization algorithm.
A brief comparison with other Python packages for Bayesian optimization showed that NUBO has competitive performance while providing a transparent and simple implementation. This makes NUBO a good candidate for optimizing expensive black-box functions when transparency is vital.
In the future, we plan to extend NUBO to include optimization strategies for multi-fidelity, multi-objective, and high-dimensional problems.
@@
The results in this paper were obtained using Python 3.11.2 with the following packages: NUBO 1.0.3, Torch 2.0.0, GPyTorch 1.10 , SciPy 1.10.1, NumPy 1.24.2, and Matplotlib 3.9.0 [40]. For the package comparison bayes_opt 1.4.3, BoTorch 0.8.4, pyGPGO 0.5.0 and SMAC3 2.0.0 were used. Python itself is available from the Python website at https://www.python.org/ and all packages used are available from the Python Package Index (PyPI) at https://www.pypi.org/.

### v3

@@ 5 Conclusion
This article introduces NUBO, a Python package for Bayesian optimisation to optimise expensive-to-evaluate black-box functions, for example, computer simulators and physical experiments. The main objective of NUBO is to make Bayesian optimisation accessible to researchers from all disciplines by providing a transparent and user-friendly implementation.
NUBO includes five sub-modules that implement Gaussian processes, acquisition functions, optimisers, test functions, and utilities. These modules provide all necessities for sequential single-point, parallel multi-point, and asynchronous optimisation of expensive-to-evaluate black-box functions for bounded, constrained, and/or mixed (discrete and continuous) input parameter spaces. We have introduced and explained each of these functionalities with individual code snippets and illustrated NUBO’s general workflow using a detailed case study that takes a hypothetical six-dimensional expensive-to-evaluate black-box function and approximates its global optimum with a parallel multi-point Bayesian optimisation algorithm.
A brief comparison with other Python packages for Bayesian optimisation showed that NUBO has competitive performance while providing a transparent and simple implementation. This makes NUBO a good candidate for the optimisation of expensive black-box functions when transparency is vital.
In the future, we plan to extend NUBO to include optimisation strategies for multi-fidelity, multi-objective, and high-dimensional problems.
@@
The results in this paper were obtained using Python 3.11.2 with the following packages: NUBO 1.0.3, Torch 2.0.0, GPyTorch 1.10 , SciPy 1.10.1, NumPy 1.24.2, and Matplotlib 3.9.0 (Hunter, 2007). For the package comparison bayes_opt 1.4.3, BoTorch 0.8.4, pyGPGO 0.5.0 and SMAC3 2.0.0 were used. Python itself is available from the Python website at https://www.python.org/ and all packages used are available from the Python Package Index (PyPI) at https://www.pypi.org/.